If a rectangle perimeter of 36ft and its 4ft wide what is its area?
Answer by RAY100(1637)

Perimeter =2l +2w = 2l +2*4 =36,,,,,,,2l = 28,,,,,l=14
.
Area = L*W = 14 *4 = 56 sq ft